# Implications of portal vector-like lepton on associated Higgs production at a multi-TeV muon collider

### v1

@@ 4 Results
The cut-flow tables corresponding to the signal events also suggest that the effect of the dark photon mass on the kinematics, in particular, on the cut efficiencies is very small as long as $m_{\gamma_{d}}<<\sqrt{s}$. However, it plays a significant role in the estimation of the cross section as it depends on values the of dark photon mass as well as the dark gauge coupling.
@@
In Fig. 15, we present the $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane at $\sqrt{s}=3$ TeV and with an integrated luminosity of 1 ab-1 in two different final states for two sets of other relevant model parameters ($g_{d},\penalty 10000\ \sin\theta_{L},\penalty 10000\ \sin\theta_{s}$).
Figure 15: $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane for two different final states at $3$ TeV muon collider center-of-mass energy for (a) $g_{d}=0.05,\penalty 10000\ \sin\theta_{L}=4\times 10^{-5},\penalty 10000\ \sin\theta_{s}=0.05$ and (b) $g_{d}=0.5,\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6},\penalty 10000\ \sin\theta_{s}=0.05$ assuming an integrated luminosity of 1 ab-1. The exclusion limits in Nc-jet = 1 (moccasin) and Nc-jet = 2 (cyan) final states are represented by dash-dot lines while the perturbative unitarity bound is represented by a dashed line (white).
@@ 5 Conclusion
A $2\sigma$ exclusion limit upto $m_{\gamma_{d}}=75\penalty 10000\ (45)$ GeV can be obtained using our analysis in the exactly one central-jet plus missing energy final state at $3$ TeV muon collider center-of-mass energy for $g_{d}=0.05\penalty 10000\ (0.5),\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}\penalty 10000\ (1.66\times 10^{-6}),\penalty 10000\ \sin\theta_{s}=0.05$ and the heavy muon mass, $m_{\mu_{p}}\geq 2$ TeV. The corresponding bounds in the exactly two central-jets plus missing energy final state are $m_{\gamma_{d}}=40\penalty 10000\ \text{GeV and}\penalty 10000\ 32$ GeV for the same set of choices of the other parameters and $m_{\mu_{p}}\geq 2.5$ TeV. Using Eq. (24) assuming $\penalty 10000\ m_{\mu_{p}}=1\penalty 10000\ \text{TeV},\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{s}=0.05$ one can exclude the value of the dark gauge coupling above $0.15$ for $m_{\gamma_{d}}\leq 90\penalty 10000\ \text{GeV}$ at $\sqrt{s}=3$ TeV with an integrated luminosity of $1$ ab-1 in the exactly one central-jet plus missing energy final state.

### v2

@@ 4 Results
The cut-flow tables corresponding to the signal events also suggest that the effect of the dark photon mass on the kinematics, in particular, on the cut efficiencies is very small as long as $m_{\gamma_{d}}\ll\sqrt{s}$. However, it plays a significant role in the estimation of the cross section as it depends on values the of dark photon mass as well as the dark gauge coupling.
@@
In Fig. 17, we present the $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane at $\sqrt{s}=3$ TeV assuming an integrated luminosity of 1 ab-1 in two different final states for two sets of $g_{d}$ and $\sin\theta_{L}$ keeping $\sin\theta_{s}$ fixed at $0.05$. The same at $\sqrt{s}=10$ TeV and an integrated luminosity of $10$ ab-1 is presented in Fig. 18. The Fig. 19 represents the $2\sigma$ exclusion limit in the $g_{d}-m_{\gamma_{d}}$ plane corresponding to $\sqrt{s}=3$ TeV and an integrated luminosity of $1$ ab-1. The regions allowed by the constraint presented in Eq. (15) and the current muon g2g-2 measurements at $1\sigma$ level are represented by the one above the dashed (black) and the solid (black) lines, respectively, in all these figures.
Figure 17: $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane for two different final states at $3$ TeV muon collider center-of-mass energy for (a) $g_{d}=0.05,\penalty 10000\ \sin\theta_{L}=4\times 10^{-5},\penalty 10000\ \sin\theta_{s}=0.05$ and (b) $g_{d}=0.5,\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6},\penalty 10000\ \sin\theta_{s}=0.05$ assuming an integrated luminosity of 1 ab-1. The exclusion limits in Nc-jet = 1 (pink) and Nc-jet = 2 (blue) final states are represented by the region below the dash-dot lines. The region above the dashed (black) line is allowed by the constraint represented in Eq. (15) while the muon g2g-2 allowed region is represented by the one above solid (black) line.
Figure 18: $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane in Nc-jet = 1 final state at $10$ TeV muon collider center-of-mass energy for (a) $g_{d}=0.05,\penalty 10000\ \sin\theta_{L}=4\times 10^{-5},\penalty 10000\ \sin\theta_{s}=0.05$ and (b) $g_{d}=0.5,\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6},\penalty 10000\ \sin\theta_{s}=0.05$ assuming an integrated luminosity of 10 ab-1. The $2\sigma$ exclusion limit is represented by dash-dot line (pink). The region above the dashed (black) line is allowed by the constraint represented in Eq. (15) while the muon g2g-2 allowed region is represented by the one above solid (black) line.
Figure 19: $2\sigma$ exclusion limit (represented by dashed-dot line) in the $g_{d}-m_{\gamma_{d}}$ plane using the data corresponding to $\sqrt{s}=3$ TeV and an integrated luminosity of $1$ ab-1 in the Nc-jet = 1 (pink) and Nc-jet = 2 (blue) final states for two different sets of $\sin\theta_{L}$ and $m_{\mu_{p}}:$ (a) $\sin\theta_{L}=4\times 10^{-5}$, $m_{\mu_{p}}=1$ TeV and (b) $\sin\theta_{L}=1.66\times 10^{-6}$, $m_{\mu_{p}}=3$ TeV. The region above the dashed (black) line is allowed by the constraint represented in Eq. (15) while the muon g2g-2 allowed region is represented by the one above solid (black) line.
@@ 5 Conclusion
[…]
@@ A.3 Vertices with Higgs:
Appendix B Squared matrix element for $\mu^{+}\mu^{-}\to h\,\gamma_{d}$
The $t$- and $u$- channel diagrams for the $\mu^{+}\mu^{-}\to h\,\gamma_{d}$ process via portal vector-like lepton mediation ($\mu_{p}$) are presented in Fig 20.
Figure 20: Feynman diagrams for the process $\mu^{+}\mu^{-}\to h\,\gamma_{d}$ mediated by the portal matter $m_{\mu_{p}}$ through the (a) $t$-channel and (b) $u$-channel.
The amplitude for the $t$-channel diagram is given by
[…]
where $c_{s}^{t},\penalty 10000\ c_{p}^{t},\penalty 10000\ c_{{}_{V}}^{t}$ and $c_{{}_{A}}^{t}$ are given by
$c_{s}^{t}$ $=$ $\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}-\dfrac{m_{\mu}}{m_{\mu_{p}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1+\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{p}^{t}$ $=$ $-\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}+\dfrac{m_{\mu}}{m_{\mu_{p}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1-\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{{}_{V}}^{t}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{p}}}\dfrac{1}{D^{2}}+1\big)$
$c_{{}_{A}}^{t}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{p}}}\dfrac{1}{D^{2}}-1\big)$
The amplitude of the $u$-channel diagram is given by
[…]
where $c_{s}^{u},\penalty 10000\ c_{p}^{u},\penalty 10000\ c_{{}_{V}}^{u}$ and $c_{{}_{A}}^{u}$ are given by
$c_{s}^{u}$ $=$ $\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}-\dfrac{m_{\mu}}{m_{\mu_{P}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1+\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{p}^{u}$ $=$ $\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}+\dfrac{m_{\mu}}{m_{\mu_{P}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1-\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{{}_{V}}^{u}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{P}}}\dfrac{1}{D^{2}}+1\big)$
$c_{{}_{A}}^{u}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{P}}}\dfrac{1}{D^{2}}-1\big)$
The spin-averaged squared amplitude, neglecting the s channel dark photon mediated and t/u channel muon mediated contribution, is
$\overline{|\mathcal{M}|^{2}}$ $=$ $\overline{|\mathcal{M}_{t}|^{2}}+\overline{|\mathcal{M}_{u}|^{2}}+2\,\operatorname{Re}\!\left(\overline{\mathcal{M}_{t}\mathcal{M}_{u}^{*}}\right)$ (B.3)
where the $t$-channel contribution in the limit $m_{\mu}\to 0$ is
$\overline{|\mathcal{M}_{t}|^{2}}$ $=$ $\frac{2}{\left[\left(k_{1}-p_{1}\right)^{2}-m^{2}_{\mu_{p}}\right]^{2}}\Bigg[A_{1}\left\{2\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)-p_{1}^{2}\left(k_{1}\cdot k_{2}\right)\right\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left(k_{1}\cdot k_{2}\right)$ (B.4)
$+\frac{1}{m_{\gamma_{d}}^{2}}\Bigg(A_{1}\Big\{2\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{2}\right)\left(p_{1}\cdot p_{2}\right)-p_{1}^{2}\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-p_{2}^{2}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)$
$+\frac{p_{1}^{2}\penalty 10000\ p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\Big\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left\{\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-\frac{p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\right\}\Bigg)\Bigg]$
Similarly, the $u$- channel contribution in the same limit reads
$\overline{|\mathcal{M}_{u}|^{2}}$ $=$ $\frac{2}{\left[\left(p_{1}-k_{2}\right)^{2}-m^{2}_{\mu_{p}}\right]^{2}}\Bigg[A_{1}\left\{2\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)-p_{1}^{2}\left(k_{1}\cdot k_{2}\right)\right\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left(k_{1}\cdot k_{2}\right)$ (B.5)
$+\frac{1}{m_{\gamma_{d}}^{2}}\Bigg(A_{1}\Big\{2\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{1}\right)\left(p_{1}\cdot p_{2}\right)-p_{1}^{2}\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-p_{2}^{2}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)$
$+\frac{p_{1}^{2}\penalty 10000\ p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\Big\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left\{\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-\frac{p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\right\}\Bigg)\Bigg]$
Here, the coefficients $A_{1}$ and $A_{2}$ are as follows
$A_{1}$ $=$ $\left(c_{{}_{V}}^{2}+c_{{}_{A}}^{2}\right)\left(c_{s}^{2}+c_{p}^{2}\right)-4c_{s}\penalty 10000\ c_{p}\penalty 10000\ c_{{}_{V}}c_{{}_{A}}$ (B.6)
$A_{2}$ $=$ $\left(c_{{}_{V}}^{2}+c_{{}_{A}}^{2}\right)\left(c_{s}^{2}+c_{p}^{2}\right)+4c_{s}\penalty 10000\ c_{p}\penalty 10000\ c_{{}_{V}}c_{{}_{A}}$
with the coupling identifications
$c_{s}=c_{s}^{t}=c_{s}^{u},\qquad c_{p}=c_{p}^{t}=-\,c_{p}^{u},\qquad c_{V}=c_{V}^{t}=c_{V}^{u},\qquad c_{A}=c_{A}^{t}=c_{A}^{u}.$
The corresponding interference term is
$\overline{\mathcal{M}_{t}\mathcal{M}_{u}^{*}}$ $=$ $\frac{2}{\left[\left(k_{1}-p_{1}\right)^{2}-m_{\mu_{p}}^{2}\right]\left[\left(p_{1}-k_{2}\right)^{2}-m_{\mu_{p}}^{2}\right]}\Bigg[2A_{3}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left(k_{1}\cdot k_{2}\right)$ (B.7)
$-\frac{1}{m_{\gamma_{d}}^{2}}\Bigg(A_{3}\Big\{p_{1}^{2}\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)+p_{2}^{2}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)-\left(p_{1}\cdot p_{2}\right)\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{1}\right)$
$-\left(p_{1}\cdot p_{2}\right)\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{2}\right)+\left(p_{1}\cdot p_{2}\right)^{2}\left(k_{1}\cdot k_{2}\right)-\frac{p_{1}^{2}\penalty 10000\ p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\Big\}$
$+\frac{A_{2}}{2}m_{\mu_{p}}^{2}\left\{p_{2}^{2}\left(k_{1}\cdot k_{2}\right)-\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)\right\}\Bigg)\Bigg]$
where the coefficient $A_{3}$ takes the form
$A_{3}$ $=$ $(c^{2}_{s}-c^{2}_{p})(c_{{}_{V}}^{2}-c_{{}_{A}}^{2})$ (B.8)
In the center-of-mass frame and in the limit $m_{h},\penalty 10000\ m_{\gamma_{d}}<<E_{{}_{\rm{CM}}}$, the $\overline{|\mathcal{M}|^{2}}$ for the pure $t$- and $u$- channels and the corresponding interference contribution take the form
$\overline{|\mathcal{M}_{t}|^{2}}$ $=$ $\frac{E_{{}_{\rm CM}}^{2}}{\left(a-b\cos\theta\right)^{2}}\Bigg[A_{1}\left(\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{h}^{2}\right)+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}+\frac{1}{2m_{\gamma_{d}}^{2}}\Bigg({A_{1}}\penalty 10000\ \Bigg\{\frac{E_{{}_{\rm CM}}^{4}}{4}\penalty 10000\ (1-\cos\theta)^{2}$ (B.9)
$-\frac{\left(m_{h}^{2}+m_{\gamma_{d}}^{2}\right)E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta+m_{h}^{2}\penalty 10000\ m_{\gamma_{d}}^{2}\Bigg\}+A_{2}\penalty 10000\ {m_{\mu_{p}}^{2}}\left\{\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{\gamma_{d}}^{2}\right\}\Bigg)\Bigg]$
$\overline{|\mathcal{M}_{u}|^{2}}$ $=$ $\frac{E_{{}_{\rm CM}}^{2}}{(a+b\cos\theta)^{2}}\Bigg[A_{1}\left(\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{h}^{2}\right)+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}+\frac{1}{2m_{\gamma_{d}}^{2}}\Bigg({A_{1}}\Bigg\{\frac{E_{{}_{\rm CM}}^{4}}{4}(1+\cos\theta)^{2}$ (B.10)
$-\frac{\left(m_{h}^{2}+m_{\gamma_{d}}^{2}\right)E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta+m_{h}^{2}m_{\gamma_{d}}^{2}\Bigg\}+\,A_{2}\,{m_{\mu_{p}}^{2}}\left\{\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{\gamma_{d}}^{2}\right\}\Bigg)\Bigg]$
$\overline{\mathcal{M}_{t}\mathcal{M}_{u}^{*}}$ $=$ $\frac{{E_{{}_{\rm CM}}^{2}}}{(a^{2}-b^{2}\cos^{2}\theta)}\Bigg[\left(A_{3}\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\right)+\frac{1}{2m_{\gamma_{d}}^{2}}\Bigg({A_{3}}\Bigg\{m^{2}_{h}m^{2}_{\gamma_{d}}$ (B.11)
$-\frac{\left(E_{{}_{\rm CM}}^{2}+m_{h}^{2}+m_{\gamma_{d}}^{2}\right){E_{{}_{\rm CM}}^{2}}}{4}\sin^{2}\theta\Bigg\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left\{\frac{E_{{}_{\rm CM}}^{2}}{8}\sin^{2}\theta-m_{\gamma_{d}}^{2}\right\}\Bigg)\Bigg]$
Here,
$a=\frac{E_{\rm CM}^{2}}{2}+m_{\mu_{p}}^{2},\qquad b=\frac{E_{\rm CM}^{2}}{2}.$